Using the matrix X = [2,2,2,2,4,6] Calculate M = ( I2 - X(XT X) "-1" XT ) and show that M is an idempotent matrix (Do not convert fractions into decimal numbers). in Assignment Expert


1
Expert's answer
2022-01-24T11:04:58-0500

"\text{Given the matrix }X=\begin{bmatrix}
 2 & 2 \\
 4 & 6
\end{bmatrix}, \text{we need to calculate }M=[I_2-X(X^TX)^{-1}X^T], \text{where } I_2\\
\text{is the } 2\times 2\text{ identity matrix.}\\
\text{Now,}\\
M=[I_2-X(X^TX)^{-1}X^T]\\
\quad=[I_2-X(X^{-1}(X^T)^{-1})X^T], \text{since }(AB)^{-1}=B^{-1}A^{-1}, \forall 
\text{ invertible matrix A and B.}\\
\quad=[I_2-X(X^{-1}(X^T)^{-1}X^T)]=[I_2-X(X^{-1})I_2]\\
\quad=[I_2-I_2]=0_2, \text{where }0_2 \text{ is the }2\times 2\text{ zero matrix.}\\
\text{Thus, }\\
M=\begin{bmatrix}
 0 & 0 \\
 0 & 0
\end{bmatrix}\\
\quad\\
\quad\\
\text{Next, to show that M is idempotent we need to show that }M^2=M.\\
M^2=\begin{bmatrix}
 0 & 0 \\
 0 & 0
\end{bmatrix}\times\begin{bmatrix}
 0 & 0 \\
 0 & 0
\end{bmatrix}=\begin{bmatrix}
 0 & 0 \\
 0 & 0
\end{bmatrix}=M"